# Frustration-free free fermions and beyond

### v1

@@ V.1 Statement
where $\hat{G}$ is the projector onto the ground space, $D(\hat{\mathcal{O}}_{\bm{x}},\hat{\mathcal{O}}_{\bm{y}}^{\prime})$ is the distance between the two operators, and $c$ and $g$ are integer constants determined by details of the Hamiltonian. We include the accurate definitions of $D(\cdot,\cdot)$, $c$ and $g$ in Appendix C. The proof of this inequality is in Ref. [38] and Ref. [17]. Although this inequality was originally derived for bosonic systems, it can be extended to fermionic systems with bosonic local Hamiltonians. Importantly, the local operators $\hat{\mathcal{O}}_{\bm{x}}$ and $\hat{\mathcal{O}}_{\bm{y}}^{\prime}$ do not need to be bosonic.
@@
A finite energy gap in the thermodynamic limit guarantees exponential decay of correlation functions in the ground state, for both bosonic and fermionic systems [39]. Therefore the correlation function in Eq. (106) in the ground state implies that the Hamiltonian is gapless.
V.2 Examples
V.2.1 Tight-binding model on checkerboard lattice
@@
[…]
asymptotically for a large $|\bm{R}|$. See Appendix C for the derivation. Combined with $\|\hat{a}^{\dagger}_{\bm{0}}|\Phi\rangle\|\approx 1/\sqrt{2}$, we find the desired power-law decay. Then the bound (107) implies a low-energy excitation, which is consistent with the quadratic band touching discussed in Sec. II.3.
V.2.2 $\eta$-pairing state
@@
VI Conclusion
In this work, we have studied frustration-free fermionic systems and established fundamental properties that characterize their low-energy excitations. For free fermionic systems, we have derived a necessary and sufficient condition for frustration-freeness. In particular, we have demonstrated that, in the presence of translation symmetry, frustration-free free fermionic systems exhibit quadratic dispersions for low-energy excitations. Furthermore, we have generalized our findings to interacting and non-translation-invariant fermionic systems by extending the Gosset-Huang inequality. Specifically, we have shown that, if the ground-state correlation function decays as a power law, the finite-size scaling of gapless excitations is $O((\log L)^{2}/L^{2})$ even in the presence of interactions and/or the absence of translation symmetry.
An important direction for future studies is to relax the strict locality assumption imposed on each local Hamiltonian in our decomposition. In this work, we have focused on strictly local decompositions of total Hamiltonians and derived the frustration-free condition and quadratic low-energy dispersions. However, Ref. [44] suggests that any gapped local Hamiltonian can be decomposed into a sum of local terms with tails decaying faster than any power law, and that the ground state of the total Hamiltonian can simultaneously be a ground state of each term in the decomposition. It remains an open question whether our frustration-free condition and results on low-energy dispersions still hold in the presence of such tails in the local Hamiltonian. We leave this intriguing investigation as a future study.
@@ B.1 BdG quasi-particle energy
which implies $U_{\mathcal{P}}\bm{\psi}^{*}$ is another eigenvector of $H_{\rm BdG}$ with eigenvalue $-E$. As eigenvectors with distinct eigenvalues ±$\pm E$, we also have
@@
$\begin{split}\hat{H}=\hat{\bm{\Upsilon}}^{\dagger}H^{(+)}\hat{\bm{\Upsilon}},\end{split}$ (135)
@@
is a positive semi-definite matrix of dimension $2N\times 2N$ and rank $N^{(+)}\leq N$.
Two comments are in order: first, when we restrict ourselves to a local term in the Hamiltonian, the above is equivalent to the main text with the replacement $2E_{i}\rightarrow\mu_{i}$. Second, the discussion applies equally well when we think of $H_{\rm BdG}$ as being the BdG Hamiltonian for the full system. In particular, when the system has translation symmetry one can first block-diagonalize in momentum blocks coupling only states at ±$\pm\bm{k}$. The discussion above holds equally well, and so the quasi-particle spectrum of $\hat{H}$ at $\bm{k}$ is given by the Fourier transform of the matrix $H^{(+)}(\bm{k})$. Importantly, $H^{(+)}$ as defined (through diagonalization of $H_{\rm BdG}$) is not a local matrix in general, and so the spectrum of $H^{(+)}(\bm{k})$ can be non-analytic, say featuring only the upper Dirac cone.
@@ B.2 Frustration-free condition
However, generally speaking, the eigenspectrum of $H_{\rm{loc}}^{(+)}$ so-defined does not coincide with the quasiparticle spectrum of $\hat{H}$. For instance, $H_{\rm{loc}}^{(+)}$ is a positive semi-definite matrix with only local terms. If the system has translation invariance, we can analyze its spectrum in the momentum space, and it follows from the Dirac trick that any touching of the eigenspectrum of $H_{\rm{loc}}^{(+)}$ with 00 must be quadratic or softer in momentum. This holds even if the true quasiparticle spectrum of $\hat{H}$ has zero-energy Dirac cones.
@@
At this step, we invoke the frustration-free condition (83), which gives $\{\hat{\psi}_{\alpha},\hat{\psi}_{\beta}\}=0$, and therefore we conclude $\{\hat{\xi}_{i}\leavevmode\nobreak\ :\leavevmode\nobreak\ i=1,\dots,N^{(+)}\}$ is a set of canonical fermions. We can therefore rewrite Eq. (138) as
@@
Appendix C Gosset–Huang inequality for fermions
C.1 Settings
Let us consider a frustration-free Hamiltonian $\hat{H}=\sum_{\bm{R}\in\Lambda}\hat{H}_{\bm{R}}$ defined on a $d$-dimensional lattice $\Lambda$. Without loss of generality, we assume all $\hat{H}_{\bm{R}}$ are positive-semidefinite and have at least one zero-eigenvalue. We consider both bosonic and fermionic systems. In the case of fermionic systems, each $\hat{H}_{\bm{R}}$ is assumed to be bosonic, which means $[\hat{H}_{\bm{R}},(-1)^{\hat{F}}]=0$ where $(-1)^{\hat{F}}$ is the fermion parity. Fermionic local terms are excluded because physical interaction terms should be local, i.e., sufficiently distant local Hamiltonians commute (not anticommute) with each other.
As the first step, we deform each local Hamiltonian $\hat{H}_{\bm{R}}$ into a projector while preserving its kernel. This operation does not alter the ground states, the presence or absence of the energy gap in the thermodynamic limit, and the order of the finite-size gap. Also, note that this operation preserves the symmetry of the local Hamiltonians. If $[\hat{H}_{\bm{R}},\hat{Q}]=0$ for some Hermitian operator $\hat{Q}$, the ground states $\{|\phi_{a}\rangle\}_{a}$ of $\hat{H}_{\bm{R}}$ can be taken to eigenvectors of $\hat{Q}$. Thus the projector $\hat{H}^{\prime}_{\bm{R}}=\hat{\mathbbm{1}}-\sum_{a}|\phi_{a}\rangle\langle\phi_{a}|$ also commute with $\hat{Q}$. In the case of fermionic systems with bosonic local Hamiltonians, the deformed local Hamiltonians are also bosonic since the deformation preserves the symmetry of fermion parity.
C.2 Interaction graph
@@
The interaction graph is defined for a decomposition $\hat{H}=\sum_{\bm{R}\in\Lambda}\hat{H}_{\bm{R}}$ of a Hamiltonian $\hat{H}$, where $\Lambda$ is the index set consisting of the lattice sites and other degrees of freedom. The interaction graph for this Hamiltonian has $|\Lambda|$ vertices labeled by $\hat{H}_{\bm{R}}$ ($\bm{R}\in\Lambda$) and there is an edge between vertex $\hat{H}_{\bm{R}}$ and vertex $\hat{H}_{\bm{R}^{\prime}}$ when $[\hat{H}_{\bm{R}},\hat{H}_{\bm{R}^{\prime}}]\neq 0$. If two vertices $\hat{H}_{\bm{R}}$ and $\hat{H}_{\bm{R}^{\prime}}$ share an edge, we say that $\hat{H}_{\bm{R}}$ is adjacent to $\hat{H}_{\bm{R}^{\prime}}$. The degree $g_{\bm{R}}$ of a vertex $\hat{H}_{\bm{R}}$ is given by the number of vertices adjacent to $\hat{H}_{\bm{R}}$. The maximum degree $g$ of the graph is defined by $g\coloneqq\max_{\bm{R}}g_{\bm{R}}$. Thus, $g$ is defined by
@@
where $\hat{H}_{\bm{R}_{0}}=\hat{H}_{\bm{R}}$, $\hat{H}_{\bm{R}_{l}}=\hat{H}_{\bm{R}^{\prime}}$, and $[\hat{H}_{\bm{R}_{i}},\hat{H}_{\bm{R}_{i+1}}]\neq 0$ for $i=0,1,\cdots,l-1$. The distance between general operators $\hat{\mathcal{O}}$ and $\hat{\mathcal{O}}^{\prime}$ is defined by
@@
Strictly speaking, $D(\hat{\mathcal{O}},\hat{\mathcal{O}}^{\prime})$ is not a distance in the mathematical sense because it does not satisfy the triangle inequality. We also note that the commutator $[\hat{\mathcal{O}},\hat{\mathcal{O}}^{\prime}]$ does not appear in the definition of the distance, and thus the nonzero commutator $[\hat{\mathcal{O}},\hat{\mathcal{O}}^{\prime}]\neq 0$ does not necessarily imply a short distance between these operators. Therefore, the distance $D(\hat{\mathcal{O}},\hat{\mathcal{O}^{\prime}})$ also works for the case that $\hat{\mathcal{O}}$ and $\hat{\mathcal{O}}^{\prime}$ are fermionic local operators.
C.3 The Gosset–Huang inequality
@@
Let $\hat{H}$ be a frustration-free Hamiltonian with the decomposition $\hat{H}=\sum_{\bm{R}\in\Lambda}\hat{H}_{\bm{R}}$ where each $\hat{H}_{\bm{R}}$ is an orthogonal projector. Let $\hat{G}$ be the orthogonal projector onto the ground space and $\epsilon$ be the spectral gap. For operators $\hat{\mathcal{O}}$ and $\hat{\mathcal{O}}^{\prime}$ , the following inequality holds.
@@
$\leq 2e^{2}\exp\left\lparen-\frac{D(\hat{\mathcal{O}},\hat{\mathcal{O}}^{\prime})-1}{c-1}\sqrt{\frac{\epsilon}{g^{2}+\epsilon}}\right\rparen.$ ) . (148)
@@
To obtain a nontrivial inequality, $\hat{\mathcal{O}}$ and $\hat{\mathcal{O}}^{\prime}$ must be local in the sense that $[\hat{\mathcal{O}},\hat{H}_{\bm{R}}]=0$ for sufficiently distant $\hat{\mathcal{O}}$ and $\hat{H}_{\bm{R}}$ (and similarly for $\hat{\mathcal{O}}^{\prime}$). However, the assumption that $\hat{\mathcal{O}}$ and $\hat{\mathcal{O}}^{\prime}$ commute is not necessary. Therefore we can use the Gosset–Huang inequality to correlation functions of fermionic operators that anticommute with each other.
The Gosset–Huang inequality provides an upper bound on the finite-size gap if there exists a power-law decaying correlation function such that
$D(\hat{\mathcal{O}},\hat{\mathcal{O}}^{\prime})\sim L,\quad\frac{|\langle\Phi|\hat{\mathcal{O}}(\hat{\mathbbm{1}}-\hat{G})\hat{\mathcal{O}}^{\prime}|\Phi\rangle|}{\|\hat{\mathcal{O}}^{\dagger}|\Phi\rangle\|\|\hat{\mathcal{O}}|\Phi\rangle\|}\geq CL^{-\Delta},$ (149)
where $C$ is a positive constant and $\Delta$ is a positive constant and $L$ is the linear dimension of the system. From the assumptions and the Gosset–Huang inequality, we have
$CL^{-\Delta}$ $\leq\frac{|\langle\Phi|\hat{\mathcal{O}}(\hat{\mathbbm{1}}-\hat{G})\hat{\mathcal{O}}^{\prime}|\Phi\rangle|}{\|\hat{\mathcal{O}}^{\dagger}|\Phi\rangle\|\|\hat{\mathcal{O}}^{\prime}|\Phi\rangle\|}\lesssim\exp\left\lparen-O(L\sqrt{\epsilon})\right\rparen.$ ) . (150)
@@
C.4 One-band model

### v2

@@ VI.1 Statement
where $\hat{G}$ is the projector onto the ground space, $D(\hat{\mathcal{O}}_{\bm{x}},\hat{\mathcal{O}}_{\bm{y}}^{\prime})$ is the distance between the two operators, and $c$ and $g$ are integer constants determined by details of the Hamiltonian. We include the accurate definitions of $D(\cdot,\cdot)$, $c$ and $g$ in Appendix D. The proof of this inequality is in Ref. [69] and Ref. [18]. Although this inequality was originally derived for bosonic systems, it can be extended to fermionic systems with bosonic local Hamiltonians. Importantly, the local operators $\hat{\mathcal{O}}_{\bm{x}}$ and $\hat{\mathcal{O}}_{\bm{y}}^{\prime}$ do not need to be bosonic.
@@
A finite energy gap in the thermodynamic limit guarantees exponential decay of correlation functions in the ground state, for both bosonic and fermionic systems [70]. Therefore the correlation function in Eq. (150) in the ground state implies that the Hamiltonian is gapless. In Appendix E, we provide two free-fermion examples of such power-law decaying correlation functions, and show that the existence of a singular band touching point between the valence and conduction bands imply power-law decaying ground-state correlation functions.
VI.2 Examples
VI.2.1 Tight-binding model on checkerboard lattice
@@
$\langle\Phi|\hat{a}_{\bm{0}}(\hat{\mathbbm{1}}-\hat{G})\hat{a}^{\dagger}_{\bm{R}}|\Phi\rangle\approx\frac{1}{2\pi}\frac{R_{y}^{\,2}-R_{x}^{\,2}}{|\bm{R}|^{4}}(-1)^{R_{x}+R_{y}}$ (152)
asymptotically for a large $|\bm{R}|$. See Appendix E for the derivation. Combined with $\|\hat{a}^{\dagger}_{\bm{0}}|\Phi\rangle\|\approx 1/\sqrt{2}$, we find the desired power-law decay. Then the bound (151) implies a low-energy excitation, which is consistent with the quadratic band touching discussed in Sec. II.3.
VI.2.2 $\eta$-pairing state
@@
VII Conclusion
In this work, we have studied frustration-free fermionic systems and established fundamental properties that characterize their low-energy excitations. For free fermionic systems, we have derived a necessary and sufficient condition for frustration-freeness. In particular, we have demonstrated that, in the presence of translation symmetry, frustration-free free fermionic systems exhibit quadratic dispersions for low-energy excitations. Furthermore, we have generalized our findings to interacting and non-translation-invariant fermionic systems by extending the Gosset-Huang inequality. Specifically, we have shown that, if the ground-state correlation function decays as a power law, the finite-size scaling of gapless excitations is $O((\log L)^{2}/L^{2})$ even in the presence of interactions and/or the absence of translation symmetry. This result implies, among other things, that the effective field theory of gapless frustration-free fermionic systems cannot be of Dirac type; in particular, it must break Lorentz invariance.
@@ B.1 BdG quasi-particle energy
which implies $U_{\mathcal{P}}\bm{\psi}^{*}$ is another eigenvector of $H_{\rm BdG}$ with eigenvalue E-E. As eigenvectors with distinct eigenvalues $\pm E$, we also have
@@
[…]
@@
is a positive semidefinite matrix of dimension $2N\times 2N$ and rank $N^{(+)}\leq N$.
Two comments are in order: first, when we restrict ourselves to a local term in the Hamiltonian, the above is equivalent to the main text with the replacement $2E_{i}\rightarrow\mu_{i}$. Second, the discussion applies equally well when we think of $H_{\rm BdG}$ as being the BdG Hamiltonian for the full system. In particular, when the system has translation symmetry one can first block-diagonalize in momentum blocks coupling only states at $\pm\bm{k}$. The discussion above holds equally well, and so the quasi-particle spectrum of $\hat{H}$ at $\bm{k}$ is given by the Fourier transform of the matrix $H^{(+)}(\bm{k})$. Importantly, $H^{(+)}$ as defined (through diagonalization of $H_{\rm BdG}$) is not a local matrix in general, and so the spectrum of $H^{(+)}(\bm{k})$ can be non-analytic, say featuring only the upper Dirac cone.
@@ B.2 Frustration-free condition
However, generally speaking, the eigenspectrum of $H_{\rm{loc}}^{(+)}$ so-defined does not coincide with the quasiparticle spectrum of $\hat{H}$. For instance, $H_{\rm{loc}}^{(+)}$ is a positive semidefinite matrix with only local terms. If the system has translation invariance, we can analyze its spectrum in the momentum space, and it follows from the Dirac trick that any touching of the eigenspectrum of $H_{\rm{loc}}^{(+)}$ with 0 must be quadratic or softer in momentum. This holds even if the true quasiparticle spectrum of $\hat{H}$ has zero-energy Dirac cones.
@@
At this step, we invoke the frustration-free condition (86), which gives $\{\hat{\psi}_{\alpha},\hat{\psi}_{\beta}\}=0$, and therefore we conclude $\{\hat{\xi}_{i}\penalty 10000\ :\penalty 10000\ i=1,\dots,N^{(+)}\}$ is a set of canonical fermions. We can therefore rewrite Eq. (182) as
@@
Appendix C Derivation of Eq. (110)
[…]
$\begin{aligned} {}_{ij}=\langle i|P_{\mathcal{W}_{f}}|j\rangle=\langle\Phi|\hat{c}^{\dagger}_{j}\hat{c}_{i}|\Phi\rangle.\end{aligned},$
where $\{|i\rangle\coloneqq\hat{c}^{\dagger}_{i}|0\rangle\}_{i}$ is a basis set of one-particle Hilbert space $\mathcal{H}_{1}$.
The $f$-orbitals can be expanded by $\{\hat{c}^{\dagger}_{i}\}_{i}$ as
$\hat{f}^{\dagger}_{\alpha}=\sum_{i}\hat{c}_{i}^{\dagger}u_{i\alpha},$ (189)
where $\sum_{i}u^{*}_{i\beta}u_{i\alpha}=\delta_{\alpha\beta}$ to be orthonormal. Therefore, we have
$P_{\mathcal{W}_{f}}=\sum_{\alpha}|f_{\alpha}\rangle\langle f_{\alpha}|=\sum_{\alpha}\sum_{i,j}u_{i\alpha}u^{*}_{j\alpha}|i\rangle\langle j|.$ (190)
We also introduce the orthonormal basis set of the orthogonal complement of $\mathcal{W}_{f}$, denoted by $\{|g_{\beta}\rangle\coloneqq\hat{g}^{\dagger}_{\beta}|0\rangle\}_{\beta}$, whose orbitals are defined by
$\hat{g}^{\dagger}_{\beta}=\sum_{i}\hat{c}_{i}^{\dagger}v_{i\beta},$ (191)
where $\sum_{i}v^{*}_{i\beta}v_{i\alpha}=\delta_{\alpha\beta}$ and $\sum_{i}u^{*}_{i\beta}v_{i\alpha}=0$. Since $\mathcal{H}_{1}=\mathcal{W}_{f}\oplus\mathcal{W}^{\perp}_{f}$, the creation operator $\hat{c}_{i}^{\dagger}$ can be represented as
$\hat{c}_{i}^{\dagger}=\sum_{\alpha}u^{*}_{i\alpha}\hat{f}^{\dagger}_{\alpha}+\sum_{\beta}v^{*}_{i\beta}\hat{g}^{\dagger}_{\beta}.$ (192)
The two-point correlation function can be calculated as
$\langle\Phi|\hat{c}^{\dagger}_{j}\hat{c}_{i}|\Phi\rangle$ $=\sum_{\alpha,\beta}u^{*}_{j\alpha}u_{i\beta}\langle\Phi|\hat{f}^{\dagger}_{\alpha}\hat{f}_{\beta}|\Phi\rangle+\sum_{\alpha,\beta}v^{*}_{j\alpha}v_{i\beta}\langle\Phi|\hat{g}^{\dagger}_{\alpha}\hat{g}_{\beta}|\Phi\rangle$ (193)
$+\sum_{\alpha,\beta}u^{*}_{j\alpha}v^{*}_{i\beta}\langle\Phi|\hat{f}^{\dagger}_{\alpha}\hat{g}_{\beta}|\Phi\rangle+\sum_{\alpha,\beta}v^{*}_{j\alpha}u^{*}_{i\beta}\langle\Phi|\hat{g}^{\dagger}_{\alpha}\hat{f}_{\beta}|\Phi\rangle.$
$=\sum_{\alpha}u_{i\alpha}u^{*}_{j\alpha},$
where we use
$\langle\Phi|\hat{f}^{\dagger}_{\alpha}\hat{f}_{\beta}|\Phi\rangle=\delta_{\alpha\beta},\langle\Phi|\hat{g}^{\dagger}_{\alpha}\hat{g}_{\beta}|\Phi\rangle=\langle\Phi|\hat{f}^{\dagger}_{\alpha}\hat{g}_{\beta}|\Phi\rangle=0.$ (194)
Therefore, we conclude
$[P_{\mathcal{W}_{f}}]_{ij}=\langle i|P_{\mathcal{W}_{f}}|j\rangle=\sum_{\alpha}u_{i\alpha}u^{*}_{j\alpha}=\langle\Phi|\hat{c}^{\dagger}_{j}\hat{c}_{i}|\Phi\rangle.$ (195)
Appendix D Gosset–Huang inequality for fermions
D.1 Settings
Let us consider a frustration-free Hamiltonian $\hat{H}=\sum_{\bm{R}\in\Lambda}\hat{H}_{\bm{R}}$ defined on a $d$-dimensional lattice $\Lambda$. Without loss of generality, we assume all $\hat{H}_{\bm{R}}$ are positive-semidefinite and have at least one zero-eigenvalue. We consider both bosonic and fermionic systems. In the case of fermionic systems, each $\hat{H}_{\bm{R}}$ is assumed to be bosonic, which means $\hat{H}_{\bm{R}},(-1)^{\hat{F}}]=0$ where $(-1)^{\hat{F}}$ is the fermion parity. Fermionic local terms are excluded because physical interaction terms should be local, i.e., sufficiently distant local Hamiltonians commute (not anticommute) with each other.
D.2 Interaction graph
@@
The interaction graph is defined for a decomposition $\hat{H}=\sum_{\bm{R}\in\Lambda}\hat{H}_{\bm{R}}$ of a Hamiltonian $\hat{H}$, where $\Lambda$ is the index set consisting of the lattice sites and other degrees of freedom. The interaction graph for this Hamiltonian has $|\Lambda|$ vertices labeled by $\hat{H}_{\bm{R}}$ ($\bm{R}\in\Lambda$) and there is an edge between vertex $\hat{H}_{\bm{R}}$ and vertex $\hat{H}_{\bm{R}^{\prime}}$ when $\hat{H}_{\bm{R}},\hat{H}_{\bm{R}^{\prime}}]\neq 0$. If two vertices $\hat{H}_{\bm{R}}$ and $\hat{H}_{\bm{R}^{\prime}}$ share an edge, we say that $\hat{H}_{\bm{R}}$ is adjacent to $\hat{H}_{\bm{R}^{\prime}}$. The degree $g_{\bm{R}}$ of a vertex $\hat{H}_{\bm{R}}$ is given by the number of vertices adjacent to $\hat{H}_{\bm{R}}$. The maximum degree $g$ of the graph is defined by $g\coloneqq\max_{\bm{R}}g_{\bm{R}}$. Thus, $g$ is defined by
@@
where $\hat{H}_{\bm{R}_{0}}=\hat{H}_{\bm{R}}$, $\hat{H}_{\bm{R}_{l}}=\hat{H}_{\bm{R}^{\prime}}$, and $\hat{H}_{\bm{R}_{i}},\hat{H}_{\bm{R}_{i+1}}]\neq 0$ for $i=0,1,\cdots,l-1$. The distance between general operators $\hat{\mathcal{O}}$ and $\hat{\mathcal{O}}^{\prime}$ is defined by
@@
Strictly speaking, $D(\hat{\mathcal{O}},\hat{\mathcal{O}}^{\prime})$ is not a distance in the mathematical sense because it does not satisfy the triangle inequality. We also note that the commutator $\hat{\mathcal{O}},\hat{\mathcal{O}}^{\prime}]$ does not appear in the definition of the distance, and thus the nonzero commutator $\hat{\mathcal{O}},\hat{\mathcal{O}}^{\prime}]\neq 0$ does not necessarily imply a short distance between these operators. Therefore, the distance $D(\hat{\mathcal{O}},\hat{\mathcal{O}^{\prime}})$ also works for the case that $\hat{\mathcal{O}}$ and $\hat{\mathcal{O}}^{\prime}$ are fermionic local operators.
D.3 The Gosset–Huang inequality
@@
Let $\hat{H}$ be a frustration-free Hamiltonian with the decomposition $\hat{H}=\sum_{\bm{R}\in\Lambda}\hat{H}_{\bm{R}}$ where each $\hat{H}_{\bm{R}}$ is a positive semidefinite operator Let $\hat{G}$ be the orthogonal projector onto the ground space and $\epsilon$ be the spectral gap. For operators $\hat{\mathcal{O}}$ and $\hat{\mathcal{O}}^{\prime}$ , the following inequality holds.
@@
[…]
@@
To obtain a nontrivial inequality, $\hat{\mathcal{O}}$ and $\hat{\mathcal{O}}^{\prime}$ must be local in the sense that $\hat{\mathcal{O}},\hat{H}_{\bm{R}}]=0$ for sufficiently distant $\hat{\mathcal{O}}$ and $\hat{H}_{\bm{R}}$ (and similarly for $\hat{\mathcal{O}}^{\prime}$). However, the assumption that $\hat{\mathcal{O}}$ and $\hat{\mathcal{O}}^{\prime}$ commute is not necessary. Therefore we can use the Gosset–Huang inequality to correlation functions of fermionic operators that anticommute with each other.
Now, suppose that there exist operators $\hat{\mathcal{O}}$, $\hat{\mathcal{O}}^{\prime}$, separated by a distance of order the system size $L$, such that their correlation function shows the power-law decay
$\frac{|\langle\Phi|\hat{\mathcal{O}}(\hat{\mathbbm{1}}-\hat{G})\hat{\mathcal{O}}^{\prime}|\Phi\rangle|}{\|\hat{\mathcal{O}}^{\dagger}|\Phi\rangle\|\|\hat{\mathcal{O}}|\Phi\rangle\|}\geq CL^{-\Delta},$ (200)
where $C$ and $\Delta$ are positive constants. Then it follows by the Gosset–Huang inequality (199) that
[…]
@@
Appendix E Power-law decaying correlation functions
In this Appendix, we provide two examples of power-law decaying correlation functions satisfying the assumption in Eq. (200). We also provide a general argument proving that singular band touching points lead to power-law decaying correlation functions.
E.1 One-band model